Supplemental Operations Report April 24th

Supplemental Operations Report 24 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Still being manually run at night from about 9 pm to 7:30 am

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 57 %

Propane Reading, director tank— 58 %

Propane Reading, intern tank— 68 %

Propane Reading, generator—43 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: no

Hours: 167.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: All rovers were serviced. The batteries were cleaned, neutralized and the cells of the batteries were refilled.

Spirit showed some corrosion and batteries were evenly depleted of water.

Curiosity showed some corrosion and batteries were unevenly depleted of water. The driver’s side had water levels equal to that of Spirit, but the passenger’s side were much more depleted. I think they were accidentally not filled the last time.

Percy showed limited corrosion and batteries were evenly depleted of water. One battery took some effort to refill due to a rusted bolt.

Opportunity showed some corrosion and batteries were evenly depleted of water.

Sojourner showed heavier corrosion and batteries were evenly depleted of water.

It took about 70 gallons of water to clean them and 13 gallons of distilled water to fill them. They will need to be serviced again during the work party.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water

CrewCar used and why, where? Yes, to Grand Junction. I fixed the broken piece on the interior driver’s side by the windshield. It just needed to pushed securely back into place.

Luna used and why, where? Yes, to Denver. Got a warning that the tire sensor is inoperable and needs to be reset. I will work on figuring this one out. Also the tires showed four different pressures right before this, so I will adjust them as needed.

Campus wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Dave and Sergii got here on Saturday and helped me do things before the arrival of the crew. This was deeply appreciated.

Summary of general operations: Nothing to report

Summary of internet: We removed the multiple sets of HudgesNet and old internet wiring throughout the Hab. We also removed the blue ethernet wire that had been put in place for the webcams that could not be used. Things look much neater as a result.

Summary of suits and radios: Dave is working with Scott on some troubleshooting of some of the suits and batteries.

Summary of Hab operations: Dave did some troubleshooting of the furnace and it is semi-operational. It can be turned on to warm the Hab but it is not completely safe as he had to bypass an exhaust sensor to do so. In addition, the new thermostat does not have an on/off switch and turns on at 50 degrees, so it will need to be disabled after each use by disconnecting the wires. This is a good enough work around for this crew, and as they are the last crew that will need it before a new furnace is installed, it will work. They will not run the furnace at night and in a few days it should be warm enough to retire the furnace forever. Dave also reset the water alarm in the loft tank.

Summary of Outpost operations: The new pump and hose have arrived for the water tank and I will set it up when I need more water.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: I dropped a 10mm wrench from the newer set between the batteries of Percy and could not retrieve it. There is an older 10 mm single wrench that can replace it until we can get it out from where it is wedged.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report


Copyright © The Mars Society. All rights reserved. | Main Site